# Does LLM Focus on the Right Words? Mitigating Context Bias in LLM-based Recommenders

### v2

@@ 5.3. Group Distributionally Robust Optimization
Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) is an optimization framework that operates over predefined sample groups, aiming to achieve consistent and reliable performance across them by emphasizing the optimizing of the worst-performing group during training. It has been widely applied in various domains (Sagawa et al., 2019; Zhao et al., 2023; Oren et al., 2019) and has demonstrated strong effectiveness in mitigating group disparities (Hashimoto et al., 2018; Zhao et al., 2023) as well as reducing models’ reliance on shortcut correlations (Han et al., 2024; Creager et al., 2021). Several studies have also integrated Group DRO into RS to improve model performance for sparse users and items (Wen et al., 2022; Zhao et al., 2023).
@@ Appendix A Appendices
A.1. Verification of Context Bias Across Diverse Prompt Templates and LLMs
In this section, we extend the FAA experiments described in Section 2.2.1 to a broader range of prompt templates (see Figure 13) and LLM backbones (see Figure 14), in order to assess whether context bias is a widespread phenomenon. For prompt templates, we adopt those proposed in (Liao et al., 2024b) and (Bao et al., 2024), referred to as Prompt1 and Prompt2. For LLM backbones, we evaluate LLaMA3‑8B (Dubey et al., 2024) and Qwen2.5‑1.5B (Team et al., 2024). Across all prompt and backbone configurations, the results consistently reveal pronounced context bias, wherein SFT leads the model to over-rely on auxiliary tokens.
@@ A.1. Verification of Context Bias Across Diverse Prompt Templates and LLMs
A.2. The proof of Lemma 3.1

### v3

@@ 5.3. Group Distributionally Robust Optimization
Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) is an optimization framework that operates over predefined sample groups, aiming to achieve consistent and reliable performance across them by emphasizing the optimizing of the worst-performing group during training. It has been widely applied in various domains (Sagawa et al., 2019; Zhao et al., 2023; Oren et al., 2019) and has demonstrated strong effectiveness in mitigating group disparities (Hashimoto et al., 2018; Zhao et al., 2023) as well as reducing models’ reliance on shortcut correlations (Han et al., 2024; Creager et al., 2021). Several studies have applied Group DRO to RS. For example, S-DRO (Wen et al., 2022) uses group DRO to improve the experience of underrepresented user groups that tend to engage with less popular items. PDRO (Zhao et al., 2023) extends this approach with popularity-aware mechanisms to prevent harming the performance of popular items.
@@ 6. Conclusion
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (62372399, 62476244), OPPO Research Fund, the Starry Night Science Fund of Zhejiang University Shanghai Institute for Advanced Study (SN-ZJU-SIAS-001), and the advanced computing resources provided by the Supercomputing Center of Hangzhou City University. We thank the reviewers for their valuable and insightful suggestions that improve the paper.
@@ Appendix A Appendices
A.1. Additional Validation of Context Bias Across Prompt Templates and LLMs
In this section, to further rule out the potential influence of prompt templates and different LLM backbones on the analysis of context bias, we extend the FAA experiments described in Section 2.2.1 to a broader range of prompt templates (see Figure 13) and LLM backbones (see Figure 14), beyond those used in the main experiments. Specifically, for prompt templates, we adopt those proposed in (Liao et al., 2024b) and (Bao et al., 2024), which we refer to as Prompt1 and Prompt2, respectively. For LLM backbones, we evaluate LLaMA3-8B (Dubey et al., 2024) and Qwen2.5-1.5B (Team et al., 2024). Across all prompt template and backbone configurations, SFT consistently amplifies the ratio of attribution values between auxiliary tokens and user-interaction tokens. This observation indicates that context bias persistently exists across different prompt templates and LLM backbones. The underlying reason is that the origin of context bias lies in the dataset itself: the co-occurrence rate between auxiliary tokens and target item tokens is significantly higher than that between interaction tokens and target item tokens, as discussed in Section 2.2.3.
@@ A.1. Additional Validation of Context Bias Across Prompt Templates and LLMs
A.2. Details of Baselines
The methods compared fall into several categories:
  • Traditional RS (SASRec (Kang and McAuley, 2018), SASRec++ (Lepage et al., 2025), DROS (Yang et al., 2023)): SASRec is a representative sequential recommendation model that employs self-attention mechanisms to effectively capture users’ dynamic interest patterns from historical interaction sequences. Building upon SASRec, SASRec++ introduces an improved training objective by adopting the softmax loss instead of BCE loss used in original SASRec, which leads to more stable optimization and enhanced recommendation performance. DROS incorporates Distributionally Robust Optimization (DRO) into sequential recommendation, aiming to improve model robustness under distributional shifts.
  • LLM-based RS (SFT (Bao et al., 2025), CFT (Zhang et al., 2024b), MSL (Wang et al., 2025b), LLaRA (Liao et al., 2024b), A-LLM (Kim et al., 2024)): This line of work leverages the strong representation and reasoning capabilities of LLMs for recommendation. SFT applies instruction-tuning strategies with carefully designed templates to adapt LLMs to recommendation tasks. CFT incorporates a causal loss to strengthen the behavior sequence modeling capabilities of LLMs. MSL improves the loss function specifically tailoring it to optimize recommendation-oriented objectives. LLaRA enhances LLM-based recommenders by incorporating embeddings from traditional recommendation models, enabling better exploitation of collaborative filtering signals. A-LLM extends this idea by aligning these collaborative embeddings with their corresponding textual semantics, facilitating more effective integration of structured and unstructured information.
  • Debiasing Methods for LLM-based RS (Reweight (Jiang et al., 2024), SPRec (Gao et al., 2025a), D3 (Bao et al., 2024)): These methods focus on mitigating various biases that arise when applying LLMs to recommendation. Reweight addresses popularity bias by balancing recommendations using pre-calculated item weights. SPRec proposes a popularity-aware negative sampling strategy within Direct Preference Optimization (DPO) (Rafailov et al., 2024) to reduce popularity bias. D3 focuses on mitigating amplification bias during inference by improving the decoding strategy, preventing the model from over-recommending items whose textual representations contain tokens with excessively high generation probabilities.
A.3. Additional Performance Comparison Across Prompt Templates and LLMs
In this section, we present additional comparative experiments on the performance of GDRT and SFT that go beyond the prompt templates and LLM backbones used in the main experiments. As summarized in Tables 4 and 5, GDRT consistently improves recommendation accuracy while achieving substantial gains in fairness across all evaluated configurations, demonstrating strong generalization ability over a broader range of prompts and LLMs.
Table 4. Performance comparison of SFT and GDRT across different prompt templates. The best result is bolded.
Prompt Dataset Method NDCG@5 DGU@5
Prompt1 (Liao et al., 2024b) Toy SFT 0.0118 0.6549
GDRT 0.0144 0.4616
Clothing SFT 0.0033 0.4631
GDRT 0.0052 0.2025
Prompt2 (Bao et al., 2024) Toy SFT 0.0116 0.5765
GDRT 0.0136 0.3987
Clothing SFT 0.0039 0.5231
GDRT 0.0045 0.1893
Table 5. Performance comparison of SFT and GDRT across different LLMs. The best result is bolded.
LLM Dataset Method NDCG@5 DGU@5
Llama3-8B Toy SFT 0.0151 0.6861
GDRT 0.0173 0.5970
Clothing SFT 0.0039 0.5801
GDRT 0.0062 0.2068
Qwen2.5-1.5B Toy SFT 0.0098 0.4849
GDRT 0.0117 0.2311
Clothing SFT 0.0018 0.3538
GDRT 0.0026 0.0778
A.4. The proof of Lemma 3.1